Session: Oct 2013 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Rubén Fernández at Thu Oct 17 07:03:33 2013) Success:

Badary (Zhukova at Thu Oct 17 07:09:24 2013) Success:

Zelenchukskaya (Kiselev P. at Thu Oct 17 07:09:58 2013) Success:

Svetloe (Andreeva T. at Thu Oct 17 07:10:56 2013) Minor failures: ANT
scans No0017-No0034 missed due to power outage, scan No0047 failed due to antenna problem

Urumqi (Wenjun Yang at Thu Oct 17 07:20:47 2013) Success:
the FS time and Mark5b time all faster 1 second than GPS time. The correlator center have used +1s to got friges. It is clear.

Medicina (scalambra at Thu Oct 17 09:15:01 2013) Success:

Noto (S. Buttaccio at Thu Oct 17 10:23:09 2013) Success:RECE
Only LCP available. Swapped polarization on K band receiver, corrected only at the end of observation.

Onsala (Michael Lindqvist at Thu Oct 17 13:03:23 2013) Success:
Bad weather rain.

Metsahovi (Elizaveta Rastorgueva-Foi at Thu Oct 17 19:02:11 2013) Failure:RECE ANT
LCP channel in RX is broken (oscillates). Wrong antcon offsets caused observations off-source

Effelsberg (Georgi at Fri Oct 18 06:45:33 2013) Success:

Torun (Pawel Wolak at Fri Oct 18 13:01:27 2013) Success:
No known major problems. Fringes to Tr and Td. LHC and RHC transposed in DBBC.

Hartebeesthoek (Jonathan Quick at Fri Oct 18 13:25:15 2013) Success:
No known problems

Shanghai (Bo Xia at Wed Oct 23 03:20:19 2013) Success: